At the surface of a freshwater lake the pressure is 106 kPa.

(a) What is the pressure increase in going 34.9 m below the surface?in kPa

(b) What is the approximate pressure decrease in going 39 m above the surface? Air at 20
Answer #1

a.)pressure increase=rho*g*h

=1000*9.81*34.9

=342.269 Kpa

b.)Decrease in pressure=1.2*9.81*39

=0.459 KPa ans
Answer #3
  1. P2-P1=?*g*h

?P=1000*9.8*34.9

?P=342.02KPa

  1. ?P=?*g*h

=1.2*39*9.8

= 458.64Pa